solve the logarithmic equation
lnx-ln(X-4)=ln3

Answer by Nate(3500)

ln(x) - ln(x-4) = ln(3) original problem
ln((x)/(x-4)) = ln(3) use quotient property of logarithms
(x)/(x-4) = 3 use the idea of ln x = ln x is also x = x
x = 3x - 12 multiply
-2x = -12 subtract
x = 6 divide